Find the exact length of the curve. x = 3 + 12t^2, y = 8 + 8t^3, 0 ≤ t ≤ 1
Answer #1

we have

x = 3 + 12t^2 \ \ and \ \ y = 8 + 8t^3

\frac{dx}{dt} =\frac{d}{dt} 3 + 12t^2 \ \ and \ \ \frac{dy}{dt} = \frac{d}{dt}8 + 8t^3

\frac{dx}{dt} =24t \ \ and \ \ \frac{dy}{dt} = 24t^2

the length of the curve,

L=\int_{a}^{b}\sqrt{\left (\frac{dx}{dt} \right )^2+\left (\frac{dy}{dt} \right )^2}dt

put a = 0, b = 1 and the value of dx/dt, dy/dt in above equation,

L=\int_{0}^{1}\sqrt{\left (24t \right )^2+\left (24t^2 \right )^2}dt

L=\int_{0}^{1}24t\sqrt{1+t^2 }dt

let assume that,

u=1+t^2

du=2tdt

and limit is,

t=0 \Rightarrow u=1+0^2=1

t=1 \Rightarrow u=1+1^2=2

we can say that,

L=\int_{1}^{2} 12\sqrt{u }du

L=\int_{1}^{2} 12u^\frac{1}{2}du

L=\left [ \frac{12u^\frac{3}{2}}{\frac{3}{2}} \right ]_{1}^{2}

L=\left [ 8u^\frac{3}{2} \right ]_{1}^{2}

L=8(2^\frac{3}{2}-1^\frac{3}{2})

L=8(2^\frac{3}{2}-1)

L=14.62741